solve each of the following systems of equations graphically if possible. indicate whether the system has a unique solution, infinitle many solutions or no solution.
3y= x-10
y= x-2

x+y=5
3x+3y= 15

Answer by Alan3354(69443)

solve each of the following systems of equations graphically if possible. indicate whether the system has a unique solution, infinitle many solutions or no solution.
3y= x-10
y= x-2
x+y=5
3x+3y= 15
------------
dl the (free) software at www.padowan.dk.com/graph/
You can graph these, see the solutions, change the colors, print them, and a lot of other things.